OpenStudy (anonymous):

y^2+4y-8x+4=0 in standard form please help me

OpenStudy (anonymous):

what standard form are they asking for

OpenStudy (anonymous):

y^2+4y-8x+4=0 y^2+4y+4= -8x (y+2)^2= -4*2*x